# Cognitive Memory in Large Language Models

### v1

@@ Abstract
This paper provides an in-depth exploration of memory mechanisms in Large Language Models (LLMs), analyzing the different types of memory and their roles within these models. Memory is a crucial element in the workflow of an intelligent agent, closely related to knowledge and profiling, yet it possesses distinct granularity and functional attributes. Despite LLMs’ capabilities in information retrieval and summarizing interactions, they lack stable and structured long-term memory. The integration of memory in AI systems is vital for providing context-rich responses, reducing hallucinations, enhancing data processing efficiency, and driving the self-evolution of AI systems. The paper begins by introducing the cognitive architecture of memory, categorizing it into sensory memory, short-term memory, and long-term memory. Sensory memory in humans captures fleeting information through senses, while in LLMs, it corresponds to input requests or prompts. Short-term memory in humans holds information briefly, whereas in LLMs, it processes inputs within the immediate context window. Long-term memory in humans includes explicit memory (episodic and semantic) and implicit memory (procedural), while LLMs implement long-term memory through external databases, vector stores, or graph structures. The text-based memory section discusses memory acquisition, management, and utilization. Memory acquisition involves selecting and compressing historical information, using methods such as memory selection and summarization. Memory management covers updating, accessing, and storing memories, as well as resolving conflicting memories. Memory utilization focuses on retrieval methods, including full-text search, SQL queries, and semantic search. The KV cache-based memory section introduces various KV selection and compression strategies. KV selection methods include regularity-based summarization, score-based approaches, and special token embeddings. KV compression techniques involve low-rank compression, KV merging, and multimodal compression. Memory management strategies include offloading memory, integrating with the operating system, and using shared attention mechanisms. The paper also explores parameter-based memory methods, such as LoRA, Test-Time Training (TTT), and Mixture of Experts (MoE). These methods transform memories into model parameters, enhancing memory efficiency and utilization. Finally, the paper discusses hidden-state-based memory, including chunk mechanisms, recurrent transformers, and the Mamba model. These approaches combine the concept of hidden states in RNNs with current methods to improve the processing and memory efficiency of long texts. Overall, this paper provides a comprehensive analysis of memory mechanisms in LLMs, highlighting their significance and offering guidance for future research directions.
@@ 1 Introduction
Memory, more precisely memories, constitutes a crucial foundational element of an agentic workflow and is intricately interwoven with knowledge and profiling. However, it deserves separate and focused attention due to its distinct granularity and functional attributes compared to ’knowledge’ and ’profile’. Profiling delineates the manner in which an agent comprehends its own identity (encompassing its character and ’avatar’), its functions (behavioral models) and its operational context (environment). Knowledge, on the other hand, supplies the factual information or learned representations that underpin decision-making processes. Memory, distinct from both, serves as the dynamic repository of experiences that integrates these elements and actively engages in decision-making. Despite decades of research, the consistent memory retention capabilities of large language models (LLMs) remain an area of ongoing exploration. Contemporary AI systems are capable of information retrieval, summarization of past interactions, and selective detail storage. However, they lack a stable and structured memory that can reliably endure over extended periods.
In the realm of artificial intelligence, the integration of memory is becoming increasingly significant. As the complexity of AI systems escalates, the incorporation of memory functions confers numerous substantial advantages upon these systems, propelling advancements across multiple dimensions.
@@
This paper will elucidate the various types of memory and their respective roles within an agentic workflow. It will also explore how these components coalesce in practical applications, clarify the mechanisms by which models in memory mode’remember’ information, and examine the transformative impact of generative AI on the very nature of memory. We first delve into the cognitive architecture, clarifying the categorization of memory and the distinctions between the ways humans and LLMs process memories. Subsequently, we analyze the current state of research on LLM memory. Finally, we introduce the future trend of LLM memory: cognitive memory.
@@ 2 Cognitive Architecture
Cognitive architecture is an important tool for studying human thinking and intelligent behavior, with a history spanning several decades. Kotseruba and Tsotsos (2020) reviews the development of cognitive architectures over the past 40 years, focusing on their research and application in simulating core human cognitive abilities, such as memory, perception, attention, and decision-making. By simulating psychological experiments, such as working memory tests and attention experiments, researchers have verified the ability of cognitive architectures to model human cognitive processes.
The Soar Cognitive Architecture Laird (2019) and the ACT-R Cognitive Architecture Ritter et al. (2019) are important works in the field of cognitive architecture. Soar is a general intelligent system that can handle various types of problems, from simple logical reasoning to complex strategic games, through its General Problem Solver (GPS) module. It also supports multiple learning mechanisms, such as rule-based learning and case-based learning. ACT-R focuses on simulating human cognitive processes, especially memory, learning, and decision-making. It studies the interaction between working memory and long-term memory through detailed memory models and can simulate human decision-making in complex tasks. These architectures have not only played an important role in artificial intelligence research but have also verified their ability to model human cognitive processes through psychological experiments. The ICARUS Cognitive Architecture Langley and Choi (2006) represents an emerging direction in cognitive architecture research. ICARUS aims to develop intelligent systems with autonomous learning and adaptation capabilities by simulating human cognitive processes. It supports autonomous learning mechanisms, can learn new knowledge through interaction with the environment, and handle complex tasks such as robot navigation and multitasking. The research on ICARUS has further promoted the application and development of cognitive architectures in the fields of artificial intelligence and cognitive psychology. In cognitive architecture, memory is divided into three types: sensor memory, short-term memory, and long-term memory.
@@ 2.1.2 Short-Term Memory
The context window defines how much past input an AI model can retain within a single exchange. This limitation is crucial in LLMs – give an AI a tiny context window, and it will forget what you said just moments ago. Expand that window, and it can maintain continuity over longer conversations, making responses more coherent and natural.
@@ 2.1.3 Long-Term Memory
In humans, long-term memory retains knowledge, experiences, and skills over time. Long-term memory is divided into explicit (declarative) memory, which is conscious and includes episodic memory (life events) and semantic memory (facts and concepts), and implicit (procedural) memory, which is unconscious and encompasses skills and learned tasks, as shown in Figure 2. For LLMs, LTM can be implemented through external databases, vector stores, or graph structures that keep relevant data available and allow the model to “recall” information in future queries.
@@ 2.2.2 LLMs Information
Large Language Models (LLMs) operate on fundamentally distinct principles compared to human memory when processing and storing information. These models are trained on extensive datasets that encompass diverse textual sources, including books, websites, and articles. During the training process, LLMs identify and learn the statistical patterns inherent in language, discerning the relationships between words and phrases. Unlike human memory, LLMs do not possess a memory system in the conventional sense; instead, they encode these linguistic patterns into a vast array of parameters—typically billions of numerical values. These parameters determine how the model predicts and generates responses based on input prompts. LLMs lack explicit memory storage akin to that of humans. When presented with a query, an LLM does not retain a record of previous interactions or the specific data used during training. Rather, it generates a response by calculating the most probable sequence of words based on the patterns learned from its training data. This process is facilitated by sophisticated algorithms, particularly the transformer architecture, which incorporates an attention mechanism. This mechanism enables the model to selectively focus on relevant segments of the input text, thereby producing coherent and contextually appropriate responses. In essence, the "memory" of LLMs is not a true memory system but rather an emergent property of their training process. They rely on the encoded patterns to generate responses, and any adaptation or learning occurs only through retraining on new data. This is a critical distinction from human memory, which evolves dynamically through continuous experiential input and is inherently adaptive in real time.
@@ 2.2.3 Parallels and Differences
Despite fundamental differences in information processing between humans and large language models (LLMs), several noteworthy parallels exist. Both systems rely heavily on pattern recognition to process and interpret data. In humans, pattern recognition is crucial for learning, enabling the recognition of faces, comprehension of language, and recall of past experiences. Similarly, LLMs excel at pattern recognition, utilizing their training data to learn the intricacies of language, predict subsequent words in a sequence, and generate coherent text. Context is another critical factor that influences both human memory and LLMs. For human memory, context enhances the effectiveness of information recall. For instance, being in the same environment where learning occurred can trigger related memories. LLMs also leverage context provided by input text to guide their responses. The transformer architecture allows LLMs to selectively attend to specific tokens (words or phrases) within the input, ensuring that the response aligns with the surrounding context.
@@ 3.1 Memory Acquisition
The memory acquisition includes three ways: memory (or called text) selection, memory summarization, and both.
@@ 3.1.2 Memory Summarization
COMEDY Chen et al. (2024b) faces the challenge of an excessively large dataset (over 500,000 data points) in extracting and generating conversation-level memories. To address this, COMEDY focuses on a subset of about 40,000 data points. GPT4-Turbo is used to extract conversation-level memories (including events, user, and robot descriptions), which are then edited by annotators to ensure the completeness and accuracy of information, creating a collection of conversation-level memories $M$. Additionally, COMEDY develops a dedicated LLM for generating conversation-level memories and filters out invalid samples to ensure data quality. In the memory compression task, COMEDY uses GPT4-Turbo to summarize all conversation-level memories $M$ from Task 1 and output the compressed memories $\hat{M}$. This process includes creating comprehensive user profiles, capturing the dynamic interactions between users and robots, and recording key events. To balance creativity and relevance, GPT4-Turbo generates three outputs at a temperature setting of 0.9. Annotators then refine and calibrate these outputs, correcting any inaccuracies or inconsistencies, and enhancing clarity and conciseness where necessary to ensure the summaries accurately reflect user profiles, interaction dynamics, and event records.
@@ 3.2 Memory Management
Memory management encompasses four aspects: memory update, memory access, data structure for memory storage, and handling contradictory memories.
@@ 3.2.1 Memory Update
The strategies for memory updating include the Least Recently Used (LRU) discard strategy, the Forgetting Curve strategy, the Previous Context Summary Update strategy, and the Integration of Old and New Knowledge strategy.
@@
where $R$ is the retention rate, $t$ is time since learning, $e$ is approximately 2.71828, and $S$ is memory strength, influenced by factors like learning depth and repetition. In Zhong et al. (2024) To simplify, the author models $S$ as a discrete value, initialized at 1 and increased by 1 each time a memory is recalled, resetting $t$ to 0. This reduces the forgetting probability, though real-life memory is more complex. MemoryBank combines these elements to create a comprehensive memory management system for LLMs, enhancing their ability to provide meaningful, personalized interactions over time and expanding possibilities for AI applications.
@@
Inspired by human memory, it is necessary to organize dynamic updates and idea evolution based on established operating principles, such as inserting new ideas, forgetting unimportant ones, and merging repetitive ones, so as to make the memory mechanism more natural and applicable. Starting with the memory cache storage architecture, Liu et al. (2023) uses a hash table to store self-generated ideas, where each hash index corresponds to a group of ideas containing similar thoughts. Within the same group, TiM supports the following operations to organize the ideas in memory: Insertion, which means storing new ideas into memory; Forgetting, which means removing unnecessary ideas from memory, such as contradictory ones; and Merging, which means combining similar ideas in memory, such as those with the same head entity.
@@ 3.2.3 Data Structure for Memory Storage
Aadhithya A et al. (2024) proposed a tree-based search. The Hierarchical Aggregate Tree (HAT) is defined as $HST=(L,M,A,\Sigma)$. Here, $L$ represents the hierarchical structure of the tree, $M$ denotes the set of nodes in the tree, $A$ is the aggregation function used to aggregate text information from child nodes to parent nodes, and $\Sigma$ also represents the set of nodes in the tree. An important characteristic of the HAT is that as you move down the hierarchy, the resolution increases, while moving from left to right results in more up-to-date information. This property enables the HAT to effectively manage and retrieve memory information in long text conversations. By using GPT to transform child nodes into parent nodes through prompts, the HAT can efficiently process and organize information.
@@
$F(\mathbf{x})=\arg\max\left([\mathbf{x}\mathbf{R};-\mathbf{x}\mathbf{R}]\right)\quad(1)$ (3)
Here, $\mathbf{R}$ is a random matrix of size $(d,b/2)$, and $b$ is the number of groups in memory. $[\mathbf{u};\mathbf{v}]$ denotes the concatenation of two vectors.

### v2

@@ Abstract
In this paper, we provide an in-depth exploration of memory mechanisms in Large Language Models (LLMs), analyzing the different types of memory and their roles within these models. Memory is a crucial element in the workflow of an intelligent agent, closely related to knowledge and profiling, yet it possesses distinct granularity and functional attributes. Despite the capabilities of LLMs in information retrieval and summarizing interactions, they lack stable and structured long-term memory. The integration of memory in AI systems is vital for providing context-rich responses, reducing hallucinations, enhancing data processing efficiency, and driving the self-evolution of AI systems. The paper begins by introducing the cognitive architecture of memory, categorizing it into sensory memory, short-term memory, and long-term memory. Sensory memory in humans captures fleeting information through the senses, while in LLMs, it corresponds to input requests or prompts. Short-term memory in humans holds information briefly, whereas, in LLMs, it processes inputs within the immediate context window. Long-term memory in humans includes explicit memory (episodic and semantic) and implicit memory (procedural), while LLMs implement long-term memory through external databases, vector stores, or graph structures. The text-based memory section discusses memory acquisition, management, and utilization. Memory acquisition involves selecting and compressing historical information, using methods such as memory selection and summarization. Memory management covers updating, accessing, and storing memories, as well as resolving conflicting memories. Memory utilization focuses on retrieval methods, including full-text search, SQL queries, and semantic search. The KV cache-based memory section introduces various KV selection and compression strategies. KV selection methods include regularity-based summarization, score-based approaches, and special token embeddings. KV compression techniques involve low-rank compression, KV merging, and multimodal compression. Memory management strategies include offloading memory, integrating with the operating system, and using shared attention mechanisms. The paper also explores parameter-based memory methods, such as LoRA, Test-Time Training (TTT), and Mixture of Experts (MoE). These methods transform memories into model parameters, enhancing memory efficiency and utilization. Finally, the paper discusses hidden-state-based memory, including chunk mechanisms, recurrent transformers, and the Mamba model. These approaches combine the concept of hidden states in RNNs with current methods to improve the processing and memory efficiency of long texts. Overall, this paper provides a comprehensive analysis of memory mechanisms in LLMs, highlighting their significance and offering guidance for future research directions.
@@ 1 Introduction
Memory, more precisely memories, constitutes a crucial foundational element of an agentic workflow and is intricately interwoven with knowledge and profiling. However, it deserves separate and focused attention due to its distinct granularity and functional attributes compared to ’knowledge’ and ’profile’. Profiling delineates how an agent comprehends its own identity (encompassing its character and ’avatar’), its functions (behavioral models), and its operational context (environment). Knowledge, on the other hand, supplies the factual information or learned representations that underpin decision-making processes. Memory, distinct from both, serves as the dynamic repository of experiences that integrates these elements and actively engages in decision-making. Despite decades of research, the consistent memory retention capabilities of large language models (LLMs) remain an area of ongoing exploration. Contemporary AI systems are capable of information retrieval, summarization of past interactions, and selective detail storage. However, they lack a stable and structured memory that can reliably endure over extended periods.
In the realm of artificial intelligence, the integration of memory is becoming increasingly significant. As the complexity of AI systems escalates, the incorporation of memory functions confers numerous substantial advantages upon these systems, propelling advancements across multiple dimensions Kseniase .
@@
This paper will elucidate the various types of memory and their respective roles within an agentic workflow. It will also explore how these components coalesce in practical applications, clarify the mechanisms by which models in memory mode remember information, and examine the transformative impact of generative AI on the very nature of memory. We first delve into the cognitive architecture, clarifying the categorization of memory and the distinctions between the ways humans and LLMs process memories. Subsequently, we analyze the current state of research on LLM memory. Finally, we introduce the future trend of LLM memory: cognitive memory.
@@ 2 Cognitive Architecture
Cognitive architecture is an important tool for studying human thinking and intelligent behavior, and it has a history spanning several decades. Kotseruba and Tsotsos (2020) reviews the development of cognitive architectures over the past 40 years, focusing on their research and application in simulating core human cognitive abilities, such as memory, perception, attention, and decision-making. By simulating psychological experiments, such as working memory tests and attention experiments, researchers have verified the ability of cognitive architectures to model human cognitive processes.
The Soar Cognitive Architecture Laird (2019) and the ACT-R Cognitive Architecture Ritter et al. (2019) are important works in the field of cognitive architecture. Soar is a general intelligent system that can handle various types of problems, from simple logical reasoning to complex strategic games, through its General Problem Solver (GPS) module. It also supports multiple learning mechanisms, such as rule-based learning and case-based learning. ACT-R focuses on simulating human cognitive processes, especially memory, learning, and decision-making. It studies the interaction between working memory and long-term memory through detailed memory models and can simulate human decision-making in complex tasks. These architectures have not only played an important role in artificial intelligence research but have also verified their ability to model human cognitive processes through psychological experiments. The ICARUS Cognitive Architecture Langley and Choi (2006) represents an emerging direction in cognitive architecture research. ICARUS aims to develop intelligent systems with autonomous learning and adaptation capabilities by simulating human cognitive processes. It supports autonomous learning mechanisms, can learn new knowledge through interaction with the environment, and can handle complex tasks such as robot navigation and multitasking. The research on ICARUS has further promoted the application and development of cognitive architectures in the fields of artificial intelligence and cognitive psychology. In cognitive architecture, memory is divided into three types: sensor memory, short-term memory, and long-term memory.
@@ 2.1 Categories of Memories
Human memory can be categorized from different perspectives. Firstly, based on the duration of memory retention, memory can be divided into iconic memory, short-term memory, and long-term memory. Iconic memory lasts for an extremely short period, typically between 0.25 and 2 seconds. It has a large capacity, but the information is not processed and fades quickly. For example, the fleeting impression we get when we see an object is a form of iconic memory. Short-term memory lasts for about one minute or less and has a limited capacity, usually around $7\pm 2$ information units. Information in short-term memory needs to be rehearsed to be maintained, such as when we repeat a phone number to remember it. Long-term memory, on the other hand, can last from over a minute to a lifetime. It has a vast capacity and stores information that has been deeply processed, like childhood experiences and learned knowledge.
The content of memory can be classified into iconic memory, logical memory, emotional memory, and procedural memory. Iconic memory involves the images of perceived objects, such as the memory of a painting or a melody, and is characterized by its intuitiveness and concreteness. Logical memory deals with the outcomes of logical thinking, such as concepts, formulas, and laws, focusing on understanding the essence and internal connections of things, like the memory of mathematical formulas and physical laws. Emotional memory is about the emotions or feelings experienced in the past, such as happiness or sadness, and carries a deep emotional hue. Procedural memory involves the images of bodily movements or actions, such as the memory of riding a bicycle or playing the piano, and is the foundation for learning motor skills. Based on the degree of conscious involvement, memory can be divided into explicit memory and implicit memory. Explicit memory refers to the conscious recall of past events or learned information. The content can be described in language, such as remembering places we have visited or knowledge we have acquired. Implicit memory, however, is the recall of past experiences without conscious involvement. It is usually related to motor skills and habits and does not require language description, like the unconscious memory of riding a bicycle.
In the cognitive structure, memory is divided into sensor memory, short-term memory, and long-term memory. We classify and introduce memory according to this standard.
@@ 2.1.2 Short-Term Memory
In the field of Natural Language Processing (NLP), the context window is a crucial concept in Large Language Models (LLMs). It refers to the range of contextual information that a model can consider when processing and generating text, typically measured in the number of tokens. For example, the context window of GPT-4o reaches 128,000 tokens, which means that it can process and generate longer and more complex texts in a single interaction.
The size of the context window has a profound impact on the model’s performance. A larger context window enables the model to maintain coherence in long conversations and better track multiple threads of dialogue. For instance, in complex multi-turn dialogues, the model can remember previous content to generate more relevant and natural responses. Moreover, a larger context window helps the model handle long documents, such as summarization and question-answering tasks. It allows the model to capture long-distance dependencies and global information within the document. A larger context window also enables the model to better understand semantic relationships in the context, thus generating more accurate and coherent text. From a technical implementation perspective, the attention mechanism is the core technology for realizing the context window. It allows the model to consider other relevant words in the text while processing a single word. The Transformer architecture and its self-attention mechanism are key to achieving a large context window. They enable the model to process all words in the text simultaneously, rather than just neighboring words. Additionally, positional encoding is used in the Transformer model to represent the position of words in the text, helping the model understand the order and relative positions of words.
@@ 2.1.3 Long-Term Memory
In humans, long-term memory retains knowledge, experiences, and skills over time. Long-term memory is divided into explicit (declarative) memory, which is conscious and includes episodic memory (life events) and semantic memory (facts and concepts), and implicit (procedural) memory, which is unconscious and encompasses skills and learning tasks, as shown in Figure 2. For LLMs, LTM can be implemented through external databases, vector stores, or graph structures that keep relevant data available and allow the model to “recall” information in future queries.
@@ 2.2.2 LLMs Information
Large Language Models (LLMs) operate on fundamentally distinct principles compared to human memory when processing and storing information. These models are trained on extensive datasets that encompass diverse textual sources, including books, websites, and articles. During the training process, LLMs identify and learn the statistical patterns inherent in language, discerning the relationships between words and phrases. Unlike human memory, LLMs do not possess a memory system in the conventional sense; instead, they encode these linguistic patterns into a vast array of parameters—typically billions of numerical values. These parameters determine how the model predicts and generates responses based on input prompts. LLMs lack explicit memory storage akin to that of humans. When presented with a query, an LLM does not retain a record of previous interactions or the specific data used during training. Rather, it generates a response by calculating the most probable sequence of words based on the patterns learned from its training data. This process is facilitated by sophisticated algorithms, particularly the transformer architecture, which incorporates an attention mechanism. This mechanism enables the model to selectively focus on relevant segments of the input text, thereby producing coherent and contextually appropriate responses. In essence, the "memory" of LLMs is not a true memory system but rather an emergent property of their training process. They rely on the encoded patterns to generate responses, and any adaptation or learning occurs only through retraining on new data. This is a critical distinction from human memory, which evolves dynamically through continuous experiential input and is inherently adaptive in real-time.
@@ 2.2.3 Parallels and Differences
Despite fundamental differences in information processing between humans and large language models (LLMs), several noteworthy parallels exist. Both systems rely heavily on pattern recognition to process and interpret data. In humans, pattern recognition is crucial for learning, enabling the recognition of faces, comprehension of language, and recall of past experiences. Similarly, LLMs excel at pattern recognition, utilizing their training data to learn the intricacies of language, predict subsequent words in a sequence, and generate coherent text. Context is another critical factor that influences both human memory and LLMs. For human memory, context enhances the effectiveness of information recall. For instance, being in the same environment where learning occurred can trigger related memories. LLMs also leverage the context provided by input text to guide their responses. The transformer architecture allows LLMs to selectively attend to specific tokens (words or phrases) within the input, ensuring that the response aligns with the surrounding context.
@@ 3.1 Memory Acquisition
Figure 3 shows the overall structure of memory acquisition. Memory acquisition includes three ways: memory (or called text) selection, memory summarization, and both.
@@ 3.1.2 Memory Summarization
COMEDY Chen et al. (2024b) faces the challenge of an excessively large dataset (over 500,000 data points) in extracting and generating conversation-level memories. To address this, COMEDY focuses on a subset of about 40,000 data points. GPT4-Turbo is used to extract conversation-level memories (including events, user, and robot descriptions), which are then edited by annotators to ensure the completeness and accuracy of the information, creating a collection of conversation-level memories $M$. Additionally, COMEDY develops a dedicated LLM for generating conversation-level memories and filters out invalid samples to ensure data quality. In the memory compression task, COMEDY uses GPT4-Turbo to summarize all conversation-level memories $M$ from Task 1 and output the compressed memories $\hat{M}$. This process includes creating comprehensive user profiles, capturing the dynamic interactions between users and robots, and recording key events. To balance creativity and relevance, GPT4-Turbo generates three outputs at a temperature setting of 0.9. Annotators then refine and calibrate these outputs, correcting any inaccuracies or inconsistencies, and enhancing clarity and conciseness where necessary to ensure the summaries accurately reflect user profiles, interaction dynamics, and event records.
@@ 3.2 Memory Management
Memory management encompasses four aspects: memory update, memory access, data structure for memory storage, and handling contradictory memories, as shown in Figure 4.
@@ 3.2.1 Memory Update
The strategies for memory updating include the Least Recently Used (LRU) discard strategy, the Forgetting Curve strategy, the Previous Context Summary Update strategy, and the Integration of Old and New Knowledge strategies.
@@
where $R$ is the retention rate, $t$ is the time since learning, $e$ is approximately 2.71828, and $S$ is memory strength, influenced by factors like learning depth and repetition. In Zhong et al. (2024) To simplify, the author models $S$ as a discrete value, initialized at 1 and increased by 1 each time a memory is recalled, resetting $t$ to 0. This reduces the forgetting probability, though real-life memory is more complex. MemoryBank combines these elements to create a comprehensive memory management system for LLMs, enhancing their ability to provide meaningful, personalized interactions over time and expanding possibilities for AI applications.
@@
Inspired by human memory, it is necessary to organize dynamic updates and idea evolution based on established operating principles, such as inserting new ideas, forgetting unimportant ones, and merging repetitive ones, so as to make the memory mechanism more natural and applicable. Starting with the memory cache storage architecture, Liu et al. (2023) uses a hash table to store self-generated ideas, where each hash index corresponds to a group of ideas containing similar thoughts. Within the same group, Tim supports the following operations to organize the ideas in memory: Insertion, which means storing new ideas in memory; Forgetting, which means removing unnecessary ideas from memory, such as contradictory ones; and Merging, which means combining similar ideas in memory, such as those with the same head entity.
@@ 3.2.3 Data Structure for Memory Storage
Aadhithya A et al. (2024) proposed a tree-based search. The Hierarchical Aggregate Tree (HAT) is defined as $HST=(L,M,A,\Sigma)$. Here, $L$ represents the hierarchical structure of the tree, $M$ denotes the set of nodes in the tree, $A$ is the aggregation function used to aggregate text information from child nodes to parent nodes, and $\Sigma$ also represents the set of nodes in the tree. An important characteristic of the HAT is that as you move down the hierarchy, the resolution increases while moving from left to right results in more up-to-date information. This property enables the HAT to effectively manage and retrieve memory information in long text conversations. By using GPT to transform child nodes into parent nodes through prompts, the HAT can efficiently process and organize information.
@@
$F(\mathbf{x})=\arg\max\left([\mathbf{x}\mathbf{R};-\mathbf{x}\mathbf{R}]\right)$ (3)
where $\mathbf{R}$ is a random matrix of size $(d,b/2)$, and $b$ is the number of groups in memory. $[\mathbf{u};\mathbf{v}]$ denotes the concatenation of two vectors.